OpenStudy (freckles):

Is there anyway we can show that 2[2*k(k+1)]+1 generates all odd squares?

OpenStudy (jtvatsim):

so like 1^2, 3^2, 5^2, and so on?

OpenStudy (freckles):

yes

OpenStudy (freckles):

i wonder if induction would work

OpenStudy (jtvatsim):

actually it simplifies rapidly

OpenStudy (jtvatsim):

It is true by definition.

ganeshie8 (ganeshie8):

Hint : Any odd square is of form 8k+1

OpenStudy (freckles):

oh wait k(k+1) is a product of consecutive integers 2|k(k+1)

OpenStudy (freckles):

so that is where you get your 2*2*2i+1

ganeshie8 (ganeshie8):

yes! that ends the proof unless you want to prove that any odd square is of form 8k+1

OpenStudy (jtvatsim):

I'm going to post a visual proof of this as well. :)

OpenStudy (freckles):

there was a question asked by @sourwing he wanted to prove that if n^2 is odd then n is odd and he wanted a direct proof so we can suppose then that n^2=2(2k(k+1))+1 so n^2=(2k+1)^2 and n=2k+1 do you think this is a good direct proof?

ganeshie8 (ganeshie8):

Looks legit to me ! also division algorithm is always there for proving these directly

OpenStudy (jtvatsim):

Interesting... that looks pretty clever actually...

OpenStudy (freckles):

division algorithm can you show me how to apply that

ganeshie8 (ganeshie8):

By division algorithm any number can be represented as 2k or 2k+1 if n is of form 2k then n^2 = (2k)^2 = 2(A) = even if n is of form 2k+1 then n^2 = (2k+1)^2 = 2(B)+1 = odd \(\blacksquare\)

ganeshie8 (ganeshie8):

I think thats not same as proving \( n^2 ~\text{is odd } \implies n ~\text{is odd}\) but you can use the contrapositive to use division algorithm its bit dumb compared to your trick

OpenStudy (jtvatsim):

Here's that visual proof I mentioned: